| Operation Summary for March 5, 2012 - March 7, 2012 |
|
Monday March 5 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, at 10:46 AM, Operators reported that Technicians valved in an air dryer at the Tevatron sector F1 service building. Operators reported at 11:18 AM that the Meson Fixed Target Critical Device Controller (CDC) had been tripping off every 15 minutes for that last couple of hours; in an attempt to resolve this problem Operators reset two Lambertson magnets (MLAM1 and 2) and Controls personnel changed out the CDC card, but to no avail. Controls personnel will change out the CDC chassis. Operations loaded protons into the Recycler at 11:48 AM. At 12:07 PM, two Linac quadrupole magnet power supplies (L:QPS213 and 215) alarmed and inhibited beam; Operators noted that QPS215’s output was noisy and lowered its output by a few amps. Operators search and secured the MTest area at 12:40 PM. A Muon person reported at 2:28 PM that the antiproton target station had been returned to service for the g-2 (g minus 2) experiment. At 3:17 PM, two Booster RF stations (BRF15 and 16) tripped off on repeated RF gap faults; they reset. Operations received the ORC (Operational Readiness Clearance) for the next Meson FTBF experiment T-1008 at 3:18 PM. Operations established beam to T-1008 3:40 PM. During the evening shift, a Linac RF station (LRF3) tripped off at 4:13 PM; it reset. Operations reported problems at 4:41 PM with sending beam to T-1008; beam was falling out in MI. A MI person investigated and found a power supply (I:HV703) load transfer switch in the wrong configuration; he power cycled the power supply. Operations began sending beam to T-1008 at 4:56 PM. At 5:25 PM, L:QPS108 tripped off; it reset. At 6:06 PM, LRF4 tripped off and would not reset; Operators called in Linac personnel who replaced a fuse. Linac resumed sending beam at 7:03 PM. At 9:16 PM, LRF1 tripped off; it reset. At 9:25 PM, a M01 LCW system tripped off and halted beam to Switchyard (SY); a pump had failed, which will be repaired in the morning. Operators reported at the end of shift that they had conducted a 750 KeV line tune. Tuesday March 6 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operators completed a search and secure of Neutrino Area NM3 and NM4 enclosures at 2:33 AM. At 3:31 AM, LRF4 tripped off for the first of two times in two minutes; it reset. Operations disabled beam to MiniBooNE at 7 AM in preparation for scheduled maintenance work. At 7 AM, Neutron Therapy Facility (NTF) personnel prepared to treat one patent later in the morning. Operators prepared the Meson M01 enclosure for access at 7:29 AM. At 7:31 AM, L:QPS104 and 108 tripped off; they both reset. During the day shift, Water Group personnel reported at 8:16 AM that the M01 LCW pumps were both bad and needed to be replaced. A Safety person reported at 8:44 AM that the Proton Area PC4 enclosure’s power had been restored, which fixed the problem that caused the area to be posted as an ODH area. Operations received the ORC for the SeaQuest (E906) experiment. Switchyard and Operations personnel began trying to establish beam to SeaQuest at 9:35 AM. Operators searched and secured the Meson MTest enclosures at 12:17 PM. At 12:57 PM, Operations reported that the two Meson M01 LCW pumps had been replaced, but one of them had no flow and the other only had a low flow; they will both needed to be replaced. Operations reported at 1:49 PM that Water Group personnel had removed the M01 pumps and began a rebuild. At 3:12 PM, LRF4 tripped off; it reset. During the evening shift, Operations established beam to Muon at 4:13 PM. Operations reported at 6:22 PM that the M01 LCW pump had been installed and Water technicians were bleeding air out of the system. A Linac Klystron RF station (KRF5) tripped off; it reset. A Neutrino Area beams person reported at 7:31 PM that he was done with beam for the night. Operations established beam to T-1008 at 7:40 PM. Operations disabled beam to Muon at 8:28 PM. Operators reported completing a 750 KeV line tune at 9:41 PM. The MI South Anode Power Supply tripped off at 11:33 PM; it reset. Operators completed a Booster Injection tune at 11:41 PM. Wednesday March 7 The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and the MTest experiment T-1008. During the midnight shift, LRF4 tripped off at 12:28 AM; it reset. A Booster bulk power supply (B:CEP400) tripped off 4:54 AM; it reset. Operations turned off all beam at 6 AM for the scheduled maintenance day. Operators began preparing Linac, Booster, MiniBooNE, and NuMI for access at 6:11 AM. Operators had MI and the Transfer Hall prepared for access by 6:14 AM. Operators had the MI-8 line prepared for access by 6:53 AM. Linac techs began working on LRF5 at 7:06 AM. Operators had N01 and NM2 ready for access by 7:21 AM. Operations Coordinator (OpsCo) Plans for Today The OpsCo wants today’s access and maintenance work to be conducted with safety in mind. |
|
More Information Comments and Suggestions |
| last modified 03/07/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |